# An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$

### v33

@@ Abstract
For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. We prove that no algorithm takes as input a non-negative integer $n$ and decides whether or not $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$$\wedge$ $(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}~(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))\wedge(\forall i,j,k\in\{0,\ldots,p\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k}))))$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
@@
(cf. [10, pp. 233–235]). A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
@@
We present an alternative proof of Theorem 1. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. The function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$, see [12]. The term "dominated" in the title of [12] means "eventually dominated". Flowchart 1 shows a semi-algorithm which computes $f(n)$ in the limit, see [12].
@@
3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers that concerns the arithmetic of $\mathbb{N}$
@@
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $\mathcal{S}$}, then $(a_{0},\ldots,a_{n})$ solves the following system of equations:
@@
No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence: every tuple $(x_{0},\ldots,x_{n})\in{\mathbb{N}}^{n+1}$ possesses an approximation in $\{0,\ldots,m\}^{n+1}$.
@@ 4 A short program in MuPAD that computes $f$ in the limit
5 An undecidable decision problem about a non-negative integer $n$ that concerns the arithmetic of $\mathbb{N}$
For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $\mathcal{S}$$\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $\mathcal{S}$} has a solution in $\{0,\ldots,b\}^{n+1}$. From [12] and Lemma 3 in [11], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $h$ in the limit.
@@
No algorithm takes as input non-negative integers $n$ and $m$ and returns the logical value of the following sentence:
@@
$(\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
@@
No algorithm takes as input a non-negative integer $n$ and returns the logical value of the following sentence:
@@
$(\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
$(\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
@@
For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations ${\mathcal{S}}\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The computability of $\beta$ is unknown.
@@ 7 A short program in MuPAD that computes $\beta$ in the limit
Let $n,m\in\mathbb{N}$. For every system of equations ${\mathcal{S}}\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $\mathcal{S}$} in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system $\widehat{\mathcal{S}}$, where
$\widehat{\mathcal{S}}=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
@@
By this and the inclusion $\widehat{\mathcal{S}}\supseteq{\mathcal{S}}$, $\widehat{\mathcal{S}}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
@@
It follows from Lemmas 4 and 5. ∎

### v34

# Two undecidable decision problems about a non-negative integer $n$ which have a short description in terms of arithmetic

@@ Abstract
We prove that the following decision problems (1) and (2) are algorithmically undecidable when $n\in\mathbb{N}$. (1) $\exists(y_{0},\ldots,y_{n})\in{\mathbb{N}}^{n+1}$ $\forall i,j,k\in\{0,\ldots,n\}$ $(((2^{\textstyle 2^{\textstyle 2^{\textstyle j}\cdot 3^{\textstyle k}}}+1$ divides $n)\Rightarrow$ $(y_{j}+1=y_{k}))$ $\wedge$ $((2^{\textstyle 2^{\textstyle 2^{\textstyle i}\cdot 3^{\textstyle j}\cdot 5^{\textstyle k+1}}}+1$ divides $n)$ $\Rightarrow$ $(y_{i}\cdot y_{j}=y_{k})))$. (2) $\exists p,q\in\mathbb{N}$ $((n=2^{p}\cdot 3^{q})$ $\wedge$ $\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}$ $\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$ $\forall i,j,k\in\{0,\ldots,p\}$ $(((x_{j}+1=x_{k})\Rightarrow(y_{j}+1=y_{k}))\wedge((x_{i}\cdot x_{j}=x_{k})\Rightarrow(y_{i}\cdot y_{j}=y_{k}))))$. For $n\in\mathbb{N}$, let $E_{n}=\{1=x_{k},~x_{i}+x_{j}=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $S$ has a solution in $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{f_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $f(n)$. For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The author proved earlier that the function $\beta:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every function $\delta:\mathbb{N}\to\mathbb{N}$ with a single-fold Diophantine representation. The computability of $\beta$ is unknown. We present a short program in MuPAD which for $n\in\mathbb{N}$ prints the sequence $\{\beta_{i}(n)\}_{i=0}^{\infty}$ of non-negative integers converging to $\beta(n)$.
@@
(cf. [11, pp. 233–235]). A computation in the limit of a function $f:\mathbb{N}\to\mathbb{N}$ is a semi-algorithm which takes as input a non-negative integer $n$ and for every $m\in\mathbb{N}$ prints a non-negative integer $\xi(n,m)$ such that $\lim\limits_{m\to\infty}\xi(n,m)=f(n)$.
@@
We present an alternative proof of Theorem 1. For $n\in\mathbb{N}$, $f(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a solution in ${\mathbb{N}}^{n+1}$, then $S$ has a solution in $\{0,\ldots,b\}^{n+1}$. The function $f:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$, see [13]. The term "dominated" in the title of [13] means "eventually dominated". Flowchart 1 shows a semi-algorithm which computes $f(n)$ in the limit, see [13].
@@
3 An undecidable decision problem about an ordered pair $(n,m)$ of non-negative integers which has a short description in terms of arithmetic
@@
Let $n,m\in\mathbb{N}$. For every system of equations $S\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ and $(a_{0},\ldots,a_{n})$ solves $S$, then $(a_{0},\ldots,a_{n})$ solves the following system of equations:
@@
No algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not
$((\forall k\in\{0,\ldots,n\}~(1=x_{k}\Rightarrow 1=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k})))$
@@ 4 A short program in MuPAD that computes $f$ in the limit
5 Two undecidable decision problems about a non-negative integer $n$ which have a short description in terms of arithmetic
Lemma 4.
([12, p. 110]). For non-negative integers, the equation $x+y=z$ is equivalent to a system which consists of equations of the forms $\beta+1=\gamma$ and $\alpha\cdot\beta=\gamma$.
For $n\in\mathbb{N}$, $h(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq$ $\{x_{j}+1=x_{k},~x_{i}\cdot x_{j}=x_{k}:~i,j,k\in\{0,\ldots,n\}\}$ has a solution in ${\mathbb{N}}^{n+1}$, then $S$ has a solution in $\{0,\ldots,b\}^{n+1}$. From Lemma 4 and [13], it follows that the function $h:\mathbb{N}\to\mathbb{N}$ is computable in the limit and eventually dominates every computable function $g:\mathbb{N}\to\mathbb{N}$. A bit shorter program in MuPAD computes $h$ in the limit.
@@
No algorithm takes as input non-negative integers $n$ and $m$ and decides whether or not
@@
$((\forall j,k\in\{0,\ldots,n\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
$(\forall i,j,k\in\{0,\ldots,n\}~(x_{i}\cdot x_{j}=x_{k}\Rightarrow y_{i}\cdot y_{j}=y_{k})))$
@@
No algorithm takes as input a non-negative integer $n$ and decides whether or not
@@
$\forall(x_{0},\ldots,x_{p})\in{\mathbb{N}}^{p+1}~\exists(y_{0},\ldots,y_{p})\in\{0,\ldots,q\}^{p+1}$
$((\forall j,k\in\{0,\ldots,p\}~(x_{j}+1=x_{k}\Rightarrow y_{j}+1=y_{k}))$$\wedge$
@@
Lemma 5.
([10, p. 7]). For every $\in\mathbb{N}, xyx\neq y$ implies that the numbers $2^{\textstyle 2^{\textstyle x}}+1$ and $2^{\textstyle 2^{\textstyle y}}+1$ are relatively prime.
Lemma 6.
([12, p. 110]). There exists a constructive algorithm that takes as input a Diophantine equation $D(x_{0},\ldots,x_{l})=0$ and returns a system $S$ of equations of the forms $y_{j}+1=y_{k}$ and $y_{i}\cdot y_{j}=y_{k}$ which is solvable in non-negative integers if and only if the equation $D(x_{0},\ldots,x_{l})=0$ is solvable in non-negative integers.
Theorem 5.
No algorithm takes as input a non-negative integer $n$ and decides whether or not
$\exists(y_{0},\ldots,y_{n})\in{\mathbb{N}}^{n+1}~\forall i,j,k\in\{0,\ldots,n\}$ (P)
$((2^{\textstyle 2^{\textstyle 2^{\textstyle j}\cdot 3^{\textstyle k}}}+1{\rm~divides~}n)\Rightarrow(y_{j}+1=y_{k}))\wedge((2^{\textstyle 2^{\textstyle 2^{\textstyle i}\cdot 3^{\textstyle j}\cdot 5^{\textstyle k+1}}}+1{\rm~divides~}n)\Rightarrow(y_{i}\cdot y_{j}=y_{k}))$
Proof.
If $n>0$, then we can compute a unique $(p,q)\in{\mathbb{N}}^{2}$ such that $n=2^{p}\cdot(2q+1)$. The decision problem (P) is algorithmically undecidable because we can obtain undecidability when $n>0$ and for every $i,j,k\in\{0,\ldots,n\}$
$(\max(j,k)>p)\Rightarrow(2^{\textstyle 2^{\textstyle 2^{\textstyle j}\cdot 3^{\textstyle k}}}+1{\rm~does~not~divide~}n)$
and
$(\max(i,j,k)>p)\Rightarrow(2^{\textstyle 2^{\textstyle 2^{\textstyle i}\cdot 3^{\textstyle j}\cdot 5^{\textstyle k+1}}}+1{\rm~does~not~divide~}n)$
In this case, by Lemma 5, for every system of equations
$S\subseteq\{y_{j}+1=y_{k},~y_{i}\cdot y_{j}=y_{k}:~i,j,k\in\{0,\ldots,p\}\}$
the problem of solvability of $S$ in non-negative integers $y_{0},\ldots,y_{p}$ is equivalent to the problem (P) for some $n=2^{p}\cdot(2q+1)$, where $n$ can be computed. Next, we apply Lemma 6 and a negative solution to Hilbert’s 10th problem. ∎
@@
For $n\in\mathbb{N}$, $\beta(n)$ denotes the smallest $b\in\mathbb{N}$ such that if a system of equations $S\subseteq E_{n}$ has a unique solution in ${\mathbb{N}}^{n+1}$, then this solution belongs to $\{0,\ldots,b\}^{n+1}$. The computability of $\beta$ is unknown.
@@ 7 A short program in MuPAD that computes $\beta$ in the limit
Let $n,m\in\mathbb{N}$. For every system of equations $S\subseteq E_{n}$, if $(a_{0},\ldots,a_{n})\in\{0,\ldots,m\}^{n+1}$ is a unique solution of $S$ in $\{0,\ldots,m\}^{n+1}$, then $(a_{0},\ldots,a_{n})$ solves the system $\widehat{S}$, where
$\widehat{S}=\{1=x_{k}:~(k\in\{0,\ldots,n\})\wedge(1=a_{k})\}\cup$
@@
By this and the inclusion $\widehat{S}\supseteq S$, $\widehat{S}$ has exactly one solution in $\{0,\ldots,m\}^{n+1}$, namely $(a_{0},\ldots,a_{n})$. ∎
@@
It follows from Lemmas 7 and 8. ∎